Case #:PM_MOD_22074  Opening date: 03/15/2022  Last update: 04/15/2022|
Terra MODIS encountered a SEU Command Processor and Format Processor reset today 03.15.22 (2022.074) at around 08:54z. As a result, MODIS entered standby mode. Though the data acquisition is going on, with some interruptions between 08:45z - 09:40z, the science quality will be degraded as the onboard calibrators (blackbody) is turned off and the onboard tables were reverted back to their at-launch values.
As a CP reset was encountered, subsequent patches need to be uploaded to the S/C after MODIS is recovered from the Standby mode. As a result of this, the recovery plan was divided into two parts, whereby in the first part, MODIS is to be transitioned from the standby to science mode and critical patches uploaded. The second part involves uploading more patches and gain tables to return the same configuration as before the anomaly. MCST and FOT are working to perform the recovery process. Recovery is expected to be in two phases:
- Bring the instrument back to science mode [074/22:21:28z]
- Upload revised tables to return to the pre-anomaly configuration [074/22:23:38z].
MODIS continues to operate normally with the FPA and instrument temperatures showing nominal values. While MCST continues to perform a detailed evaluation of the radiometric performance of the instrument since the anomaly, the data since the anomaly is expected to be of a degraded science quality. In particular, band 5 gains (reset to at-launch value due to this anomaly) is expected to show about 20-30% difference relative to pre-anomaly levels. We expect these to return to pre-anomaly levels once the entire recovery process is complete. We will keep you posted on the instrument recovery status and data anomaly. Further details may be available in the NRT note.
SEU Reset: UPDATE 03.15.22********************
The first part of the recovery process was completed this evening (03.15.2022) and MODIS successfully transitioned to science mode. The second part of the recovery process is to be reviewed with ART on 03.16 and subsequent recovery is planned thereafter.
SEU Reset: UPDATE 03.17.22********************
The MODIS IOT performed the second part of the commanding to upload additional patches and gain tables to successfully return the MODIS instrument to the same configuration as before the anomaly. All the commanding was successfully performed. Band 5 gains have returned to pre-anomaly levels, however, we expect degraded science data quality through about 076/02:20, when the SCI_ABNORM_FLAG was set back. A more detailed timeline of the outages is being compiled. Over the next few days, radiometric performance will continue to be evaluated and any updates will be reported.
SEU Reset: UPDATE 03.29.22********************
MODAPS will re-process all affected Terra L1B, L2, and L3 dailies and 8-day products for C6 (2022074 - 2022076) and C61 (2022074 - 2022084). This has become necessary in light of the post-event assessment performed by the MODIS Calibration Team (MCST) which suggested the following after-effects:
1) Minor gain changes that were noticed in the RSB bands, with band 5 showing significant gain change.
2) Amongst the TEB bands, band 21 is showing slightly higher noise levels.
3) Changes noticed in mirror side differences.
For more details, users may refer to the MCST findings.
SEU Reset: UPDATE 04.15.22********************
All affected Terra products, both dailies and n-days, have been re-processed for both C6 and C61. This issue is now considered resolved.
Occurrence: TERRA C61 L1B and all Downstream products DOY 2022.074 and forward.